OpenStudy (anonymous):

Find the value of csc theta , if cos theta= -3/5; 90 < theta < 180.

terenzreignz (terenzreignz):

Keeping this in mind... \[\Large \csc(\theta) = \frac1{\sin(\theta)}\]

terenzreignz (terenzreignz):

Can you try to find \(\sin(\theta)\) using the given information?

OpenStudy (anonymous):

yes thank you